How many liters of ammonia gas is generated with 4.50 L of nitrogen gas and excess hydrogen gas?
1
Expert's answer
2020-03-17T09:19:48-0400

The formation of ammonia is described by the following reaction:

N2 + 3H2 = 2NH3

As 1 mole of any gas occupies 22.4 L at standard conditions, than:

V(NH3) = V (N2) × (2 × 22.4 L) / 22.4 = 4.50 L × (2 × 22.4 L) / 22.4 L = 9.0 L


Answer: 9.0 L of ammonia